Ferrari:

Which is an equivalent expression for three sevenths squared times two thirds raised to the power of negative three all raised to the power of negative one?

Ferrari:

1 attachment
Ferrari:

seven thirds squared times two thirds cubed seven thirds to the power of negative two times two thirds to the power of negative three three sevenths squared times two thirds raised to the power of negative three three sevenths times two thirds raised to the power of negative four

Ferrari:

is it 1.61316872

Ferrari:

the expression in the question is it 1.61316872

Ferrari:

bro help

Ferrari:

nvm man i can't do this anymore

surjithayer:

\[\left[ \left( \frac{ 3 }{ 7 } \right)^2*\left( \frac{ 2 }{ 3 } \right) ^{-3}\right]^{-1}\] \[=\left[ \left( \frac{ 3 }{ 7 } \right) ^2*\left( \frac{ 3 }{ 2 } \right)^3\right]^{-1}\] \[=\left[ \frac{ 9 }{ 49 } *\frac{ 27 }{ 8 }\right]^{-1}\] \[=\left[ \frac{ 243 }{ 392 } \right]^{-1}\] \[=\frac{ 392 }{ 243 }\]